# Unreal Robotics Lab: A High-Fidelity Robotics Simulator with Advanced Physics and Rendering

### v1

@@ Abstract
High-fidelity simulation is essential for robotics research, enabling safe and efficient testing of perception, control, and navigation algorithms. However, achieving both photorealistic rendering and accurate physics modeling remains a challenge. This paper presents a novel simulation framework–the Unreal Robotics Lab (URL) that integrates the Unreal Engine’s advanced rendering capabilities with MuJoCo’s high-precision physics simulation. Our approach enables realistic robotic perception while maintaining accurate physical interactions, facilitating benchmarking and dataset generation for vision-based robotics applications. The system supports complex environmental effects, such as smoke, fire, and water dynamics, which are critical for evaluating robotic performance under adverse conditions. We benchmark visual navigation and SLAM methods within our framework, demonstrating its utility for testing real-world robustness in controlled yet diverse scenarios. By bridging the gap between physics accuracy and photorealistic rendering, our framework provides a powerful tool for advancing robotics research and sim-to-real transfer.
@@ I INTRODUCTION
Simulation plays a crucial role in robotics research, enabling safe, efficient, and cost-effective development, testing, and validation of robotic systems before real-world deployment. High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality. However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [1].
This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [2] with the advanced physics simulation of MuJoCo [3] to create photorealistic environments for robotics applications. By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects—such as smoke, fire, and water waves—which are crucial for training robots to handle diverse and unpredictable conditions. These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing simulators. As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
We present a novel, open-source simulation framework - the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation. Unlike existing simulators focused on specific domains (e.g., autonomous driving or aerial robotics) or embodied-ai for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more. Additionally, we incorporate CoACD-based convex decomposition [4] for improved non-convex object representation. Our key contributions are:
@@
  • Scene Generation: A GPT-based module automates diverse environment creation, reducing manual setup and enabling adaptive scene variations.
  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with full Robot Operating System (ROS) compatibility.
@@
The subsequent sections are structured as follows. Sec. II provides an overview of the relevant literature, discussing previous works in the field, while Sec. III outlines the proposed method in detail. Sec. IV presents the experimental setup, including both simulation and real robot experiments. Finally, in Sec. V, we present the conclusions and discuss future
@@ II RELATED WORK
Several other game engines and simulation platforms offer photorealistic rendering [5, 6, 7], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry. Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [8] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change.
Furthermore, Nanite’s automatic level of detail (LOD) [9] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance. This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical. Beyond rendering, Unreal Engine offers a flexible and extensible ecosystem, supporting third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment, making it highly adaptable for a variety of robotic simulation tasks. These features ensure that simulations can achieve both high physical accuracy and photorealistic perception, making them an ideal choice for vision-based robotics research.
@@ II-A Robotics-Focused Simulation Platforms
Robotic simulation has made considerable progress, with numerous frameworks addressing challenges in physics accuracy, rendering quality, and real-time performance. This section reviews related work in three primary areas: robotics simulation platforms, comparisons of physics engines, and benchmarking of SLAM and navigation under adverse visual conditions.
Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency. Gazebo [10] is a widely-used open-source simulator that supports multiple physics engines; ODE [10], Simbody [11], (Py)Bullet [12], and DART [13], although its rendering capabilities lag behind modern game-engine-based simulators. MuJoCo [3] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering. GPU-accelerated platforms such as Isaac Orbit [5] and Isaac Gym [14] enable large-scale reinforcement learning, prioritizing speed over advanced rendering. Intel SPEAR [15] provides photorealistic indoor environments suitable for embodied AI research. CARLA [7] and AirSim [16] leverage Unreal Engine for photorealistic urban and aerial simulations, with a primary focus on autonomous driving and drone applications. CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding. However, both simulators remain largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios. RoboSuite [17] and RoboCasa [18] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity. MetaUrban [19] simulates realistic urban environments with dynamic traffic, beneficial for multi-agent interactions.
Although most simulators prioritize either physics accuracy or visual realism, few achieve both. Our approach bridges this gap by integrating MuJoCo for precise physics with Unreal Engine for state-of-the-art rendering, creating a high-fidelity simulation pipeline for diverse robotic applications.
@@ II-B Physics Engines for Robotics Simulation
Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions. Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [20]. MuJoCo [3] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion. However, it lacks built-in high-fidelity rendering. ODE (Open Dynamics Engine) [10], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning. Bullet [12] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
Unreal Engine’s Chaos and Havok [21] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations. However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics. NVIDIA PhysX [22] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control.
Our approach combines MuJoCo’s precise physics with Unreal Engine’s advanced rendering, achieving both photorealism and accurate simulation for robotic applications.
@@ II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting.
Envodat [23] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity. Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings.
Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim). A recent simulator, SPEAR [15], integrates MuJoCo with Unreal Engine but primarily focuses on providing a Python-based OpenAI Gym environment, pre-built scenes, limited robot support, and task-specific benchmarks. In contrast, our approach integrates MuJoCo’s precise physics with Unreal Engine’s advanced rendering in a flexible and generalized simulation environment. Our framework allows users to freely configure robots and tasks, and utilizes ROS to treat simulated interactions as closely as possible to real-world scenarios. Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios.
@@ III-A Overview
Our simulator (URL) integrates MuJoCo and Unreal Engine to provide a comprehensive simulation environment that closely mimics real-world interactions. Robots can be effortlessly integrated by adding corresponding MuJoCo XML descriptions and associated Unreal Engine assets. The system supports standard robot configurations and sensors, including cameras, IMUs, and force sensors, ensuring wide compatibility. The overall system can be seen in Figure 2. The system is comprised of both External components and the internal simulator. The system has a SimManager that is modular and manages submodules comprised of a Benchmark, Replay, and SceneGen modules. It manages all references to the UE data and is developed as a plugin for UE that can be added to any project.
For communication, our simulator leverages ROS to enable seamless interaction with existing robotics frameworks. Furthermore, faster custom communication methods are being explored through ZeroMQ [24] for improved real-time performance. We treat the simulation environment as a surrogate real-world, facilitating the testing and validation of robotic behaviors and perception systems in realistic scenarios.
III-B Physics Manager
Our system integrates MuJoCo and CoACD [4] within the Unreal Engine (UE) for accurate physics simulation and mesh decomposition. These systems are currently compiled into the UE project, and extend a custom interface. To represent objects in the scene, we introduce an intermediate representation, termed PhysicsObject, which encapsulates both geometric and state information. This allows us to efficiently manage and transform scene elements between UE and MuJoCo while ensuring that transformations and dynamic interactions are correctly synchronized.
Objects within the scene are categorized into predefined types. Specifically, assets are tagged under one of the following categories: EP_STATIC, EP_STATIC_COMPLEX, EP_DYNAMIC, EP_DYNAMIC_COMPLEX, and EP_PRIMITIVE_(TYPE). Static objects remain unchanged throughout the simulation, whereas dynamic objects are actively tracked, with their transformations updated asynchronously within UE. The distinction between simple and complex dynamic objects determines whether the object undergoes convex decomposition by CoACD. CoACD decomposes meshes into a series of convex hulls, whereas the regular mujoco compile will generate only one convex hull per object. This allows CoACD to represent non-convex geometry and is recommended by MuJoCo [20]. This classification is recorded within the corresponding PhysicsObject to guide further processing.
@@ III-B Physics Manager
Once asset classification is complete, mesh-based assets are exported as OBJ files, ensuring compatibility with MuJoCo. Additionally, if landscapes are present in the scene, they are scanned and converted into height-fields to be used within the simulation. This process uses a custom collision channel on the landscape, then, utilizing an autogenerated bounding box, we perform multithreaded raycasts to determine the heights to use. This bounding-box limits the raycasts to within this area only at a predefined resolution. This prevents the whole landscape from being scanned. Figure 1 demonstrates the output of this process. This method was chosen because landscapes can be very large, and the user may not be interested in converting the entire landscape for physics. It also has the added benefit of being able to model small clutter and decorations in the heightfield itself. This process enables the seamless reconstruction of large-scale terrains within MuJoCo, preserving critical environmental features necessary for accurate physics interactions.
@@
Following asset processing, an XML scene description is automatically generated to initialize MuJoCo. This XML encapsulates all relevant object data, including mesh file references, physics properties, and hierarchical relationships. Special attention is paid to robotic assets that are detected within the MuJoCo environment and mapped back to their corresponding UE entities. During this process, cameras and sensors specified in the XML are instantiated within UE to ensure consistency across both environments. Sensor elements are represented as sites within MuJoCo, allowing for precise placement and interaction within the physics simulation.
@@
The physics simulation operates asynchronously, running within a separate execution thread to maximize performance. Transformation updates are also managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo. External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment. UE physics can still be used for objects not managed by MuJoCo.
To improve computational efficiency, assets that remain unchanged between simulation runs are cached, preventing unnecessary recomputation and redundant file exports. This caching mechanism ensures that static elements do not incur additional processing overhead, allowing the system to focus on dynamic entities that require continuous state updates. The final XML scene description, along with the automatically generated mesh assets, enables direct execution within MuJoCo as a standalone simulation. A supporting script is provided to facilitate this process, allowing the same environment to be used independently of UE when required.
III-C Benchmark System
The benchmarking system automates the recording and management of evaluation metrics throughout the simulation. At its core, it consists of a MetricManager and a BaseMetric class, providing a structured and extensible framework to track performance indicators.
The BaseMetric class defines a fundamental interface for all metrics, encapsulating a two-dimensional float array that serves as a storage mechanism for time-series data. This class provides essential methods, including initialization, resetting, and exporting data to CSV format. These methods ensure that each metric implementation adheres to a consistent lifecycle, simplifying the process of integrating new metrics into the system. Derived metrics can override these base methods to implement domain-specific logic while still benefiting from the structured data storage and export mechanisms.
The MetricManager is responsible for the automatic discovery and registration of all metric implementations. It dynamically populates a list of available metrics by scanning for corresponding class definitions, allowing modular and scalable metric tracking. Users can enable or disable specific metrics through a dedicated user interface, ensuring flexibility in selecting relevant performance indicators for a given experiment.
By default, the system tracks core metrics, including: DistanceToGoal, which tracks the remaining distance from an entity to a predefined target location; Collisions, which records contact events with obstacles or other entities; TimeToGoal, which measures the time required to reach the destination; and GlobalPose, which logs the absolute position and orientation of an entity over time.
All metrics are recorded at discrete intervals and automatically saved as CSV files within designated experiment folders, ensuring organized and accessible results. The system also features grace periods (delaying data recording for $n$ seconds) and timeouts (terminating recordings after $n$ seconds), improving efficiency and reproducibility across simulation runs.
III-D Scene Generation
Figure 3: Example of room generated using the experimental Scene Generation system leveraging GPT API.
Our simulator features an experimental scene generation module designed primarily for indoor environments, leveraging OpenAI’s ChatGPT [25] API. When an instance of the ASceneGenerator Actor is present and assigned a RoomType, the system automatically generates object placements based on the specified scene type (Fig. 3). The module queries ChatGPT with a structured prompt containing a list of available assets and receives a JSON-formatted response specifying object positions. These placements are then parsed, and the corresponding assets are instantiated in the environment.
To ensure spatial coherence, the system detects overlapping objects using Unreal Engine’s physics system. If overlaps occur, an iterative refinement loop provides feedback to ChatGPT for updated placement recommendations, continuing until all collisions are resolved or a predefined iteration limit is reached. This method enables efficient, collision-free scene generation, producing diverse and structured environments suitable for robotic experimentation.
III-E Replay

### v2

@@ Abstract
High-fidelity simulation is essential for robotics research, enabling safe and efficient testing of perception, control, and navigation algorithms. However, achieving both photorealistic rendering and accurate physics modeling remains a challenge. This paper presents a novel simulation framework, the Unreal Robotics Lab (URL), that integrates the advanced rendering capabilities of the Unreal Engine with MuJoCo’s high-precision physics simulation. Our approach enables realistic robotic perception while maintaining accurate physical interactions, facilitating benchmarking and dataset generation for vision-based robotics applications. The system supports complex environmental effects, such as smoke, fire, and water dynamics, which are critical to evaluating robotic performance under adverse conditions. We benchmark visual navigation and SLAM methods within our framework, demonstrating its utility for testing real-world robustness in controlled yet diverse scenarios. By bridging the gap between physics accuracy and photorealistic rendering, our framework provides a powerful tool for advancing robotics research and sim-to-real transfer. Our open-source framework is available at https://unrealroboticslab.github.io/.
@@ I INTRODUCTION
Simulation plays a crucial role in robotics research, enabling safe, efficient, and cost-effective development, testing, and validation of robotic systems before real-world deployment. High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality. However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [szot2021habitat]. This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [unrealengine] with the advanced physics simulation of MuJoCo [todorov2012mujoco] to create photorealistic environments for robotics applications. By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects — such as smoke, fire, and water waves — which are crucial to training robots to handle diverse and unpredictable conditions. These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing general-purpose robotics simulators. As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
We present a novel, open-source simulation framework — the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation. Unlike existing simulators focused on specific domains (e.g. autonomous driving or aerial robotics) or embodied AI for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more. While powerful general-purpose simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, our framework prioritizes physical fidelity by utilizing the MuJoCo physics engine, which is highly validated for its accuracy and stability in contact-rich control tasks. Additionally, we incorporate CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation. Our key contributions are:
@@
  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with Robot Operating System (ROS) compatibility.
@@ II RELATED WORK
Several other game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry. Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [unrealengine] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change. Furthermore, Nanite’s automatic level of detail (LOD) [unrealengine] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance. This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical. Beyond rendering, Unreal Engine offers a flexible and extensible ecosystem, supporting third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment, making it highly adaptable for a variety of robotic simulation tasks. These features ensure that simulations can achieve both high physical accuracy and photorealistic perception, making them an ideal choice for vision-based robotics research.
@@ II-A Robotics-Focused Simulation Platforms
Robotic simulation has made considerable progress, with numerous frameworks addressing challenges in physics accuracy, rendering quality, and real-time performance. This section reviews related work in three primary areas: robotics simulation platforms, comparisons of physics engines, and benchmarking of SLAM and navigation under adverse visual conditions. Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency. Gazebo [koenig2004design] is a widely-used open-source simulator that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], (Py)Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering capabilities lag behind modern game-engine-based simulators. MuJoCo [todorov2012mujoco] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering. GPU-accelerated platforms such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] enable large-scale reinforcement learning, prioritizing speed over advanced rendering. Isaac Sim [NVIDIA_Isaac_Sim], built on NVIDIA Omniverse, provides extensible robotics simulation with high-fidelity physics via PhysX [nvidia_physx]—which excels in angular stability and scalability but may lag in accuracy for contact dynamics and linear stability compared to MuJoCo. Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments suitable for embodied AI research. CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for photorealistic urban and aerial simulations, with a focus on autonomous driving and drone applications. CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding. However, both simulators are largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios. RoboSuite [robosuite2020] and RoboCasa [robocasa2024] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity. Although most simulators prioritize either physics accuracy or visual realism, few achieve both. Our approach bridges this gap by integrating MuJoCo (precise physics) with Unreal Engine (state-of-the-art rendering), creating a high-fidelity simulation pipeline for diverse robotic applications.
@@ II-B Physics Engines for Robotics Simulation
Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions. Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [erez2015simulation]. MuJoCo [todorov2012mujoco] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion. However, it lacks built-in high-fidelity rendering. ODE (Open Dynamics Engine) [koenig2004design], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning. Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
Unreal Engine’s Chaos and Havok physics solvers [havok_physics] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations. However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics. NVIDIA PhysX [nvidia_physx] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control. Our approach combines MuJoCo’s precise physics, which are validated for their accuracy and stability [erez2015simulation], with Unreal Engine’s advanced rendering. This integration achieves both photorealism and accurate simulation for robotic applications. It also provides a broad feature set by inheriting the built-in capabilities of both Unreal Engine and MuJoCo.
@@ II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting. Envodat [nwankwo2024envodat] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity. Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings. Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim). While our work shares a similar goal with recent simulators such as SPEAR [simulator_intelspear] simulator in combining Unreal Engine with MuJoCo, our frameworks are built on fundamentally different design philosophies. SPEAR is primarily structured as a Python-based OpenAI Gym environment, making it well-suited for task-specific, algorithmic research in embodied AI. In contrast, our framework is designed for systems-level validation through its native ROS integration, allowing researchers to test a complete robotics software stack as it would be deployed in the real world. This focus on flexible, real-world system replication is further reflected in our integrated workflow. Our framework allows users to build and simulate scenes directly within the Unreal Editor, automatically handling all asset processing. This streamlined approach eliminates the need for separate scripts and manual data handling for importing and exporting assets, which can be required in other simulators. Ultimately, this design makes URL a generalized platform where users can freely import any MuJoCo-compatible robot and construct custom tasks, rather than focusing on a curated set of pre-built scenes and benchmarks. Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios. Beyond rendering, UE’s mature ecosystem accelerates robotics research. We utilize its built-in NavMesh and Behavior Tree tools to script complex, dynamic scenarios with moving NPCs and vehicles, creating environments that are not just visually realistic but also behaviorally rich. This is critical for testing planner robustness in non-static worlds, a feature often lacking in physics-focused simulators.
@@ III-A Overview
Our simulator is implemented as a UE plugin centered around a custom SimManager Actor, with the overall architecture shown in Fig. 2. This Actor manages several core sub-systems: a SceneManager for processing assets and constructing the MuJoCo scene, a PhysicsManager to synchronize MuJoCo physics with UE rendering, and Benchmark and Replay systems for performance evaluation and data logging. The pipeline begins with the SceneManager scanning for actors with a UCustomPhysicsComponent, processing their properties to construct a MuJoCo Specification (MjSpec), and compiling the final physics model. The system supports standard robots defined via MuJoCo XML with corresponding UE assets, including sensors such as cameras, IMUs, and force sensors. Communication leverages a modified ROS plugin [mania19scenarios] for compatibility with existing robotics frameworks, allowing the simulator to act as a high-fidelity, real-world surrogate for testing robotics behaviors and perception. To complement the physics and data pipeline, our framework uses the larger set of features of Unreal Engine to build rich and dynamic worlds. This includes high-fidelity rendering for perception research (Lumen, Nanite), rapid world-building with automated NavMesh generation for pathfinding, and complex scenario creation using Behavior Trees for intelligent agents and the Niagara system for adverse weather and effects.
III-B Scene Building and Asset Processing
UE actors are configured for MuJoCo using a UCustomPhysicsComponent, which contains an intermediate PhysicsObject to synchronize geometry and state between the two engines. The settings of this component dictate how the object is processed for simulation. Its geometric complexity setting (Complex or Simple) determines how meshes are converted into convex hulls for MuJoCo’s collision system, using either CoACD [wei2022coacd] for decomposition or direct compiler approximation. The component also controls the object’s mobility (static or dynamic) via an override or by defaulting to the actor’s native UE setting.
@@ III-B Scene Building and Asset Processing
The SceneManger then iterates over all actors with the UCustomPhysicsComponent and processes the assets with their containing parameters. Meshes that are processed with CoACD are saved in the project’s SavedDir for caching. UE landscapes are converted to MuJoCo height fields using multithreaded raycasts within a user-defined bounding box at a set resolution, as shown in Fig. 1. This targeted approach avoids processing entire large-scale terrains while still capturing fine details such as small clutter, preserving key environmental features for the physics simulation.
@@
MjSpec is then used to generate the final MuJoCo Model and XML file. Robotic assets, packaged as UE pawns with their own XML, are automatically detected and added into the MjSpec. Sensors are defined as sites in the robot’s XML, and we employ a hybrid sensing approach: non-visual data (IMUs, force sensors) is sourced directly from MuJoCo, while visual data (cameras, LiDAR) is rendered in UE for photorealism. The exported XML and assets can also be used to run the scene as a standalone MuJoCo simulation, independent of UE.
III-C Physics Manager
@@
The physics simulation operates asynchronously, running within a separate execution thread to maximize performance. Specifically, the MuJoCo physics state is advanced by calling mj_step within a dedicated thread that maintains a configurable physics timestep (e.g., 1000 Hz) by yielding control between steps. Meanwhile, the Unreal Engine rendering thread operates independently at its own rate, typically 60 FPS or higher, depending on scene complexity. Transformation updates are managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo. To ensure data integrity, locking mechanisms are used when the rendering thread accesses the latest physics state, a design that mimics the behavior of MuJoCo’s native simulate GUI and decouples rendering performance from physics fidelity. External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment. UE physics can still be used for objects not managed by MuJoCo.
III-D Communication
Our framework uses a modified ROS plugin [mania19scenarios] that automates the communication layer. This plugin automatically publishes all sensor data from both MuJoCo and UE to their corresponding ROS topics, the names of which are derived directly from the sensor definitions in the robot’s XML file. Similarly, for robot control, the plugin dynamically generates subscribers for any defined actuators or joints. This allows standard robotics controllers to send actuator or joint commands directly to MuJoCo.
III-E Benchmark System
The benchmarking system automates metric recording and management through a structured, extensible framework built upon a MetricManager and a BaseMetric class. The BaseMetric class provides a fundamental interface for all metrics, handling time-series data storage and defining a consistent lifecycle with methods for initialization, reset, and CSV export. New metrics can be created by deriving from this class to implement domain-specific logic while inheriting the core data-handling functionality. The MetricManager automatically discovers and registers all metric classes by scanning project definitions, allowing users to select and configure them for each experiment via a UI. By default, the system tracks core metrics such as DistanceToGoal (distance to target), Collisions (contact events), TimeToGoal (task completion time), and GlobalPose (absolute pose). All metrics are recorded at discrete intervals and saved as CSV files in designated experiment folders, with support for configurable grace periods and timeouts to ensure reproducible results.
III-F Replay System